# The Existence of MacWilliams-Type Identities for the Lee, Homogeneous and Subfield Metric

### v2

@@ Acknowledgments
Jessica Bariffi was partly supported by the Federal Ministry of Education and Research of Germany in the program of "Souverän. Digital. Vernetzt." Joint project 6G-RIC, project identification number: 16KISK022. She is now funded by the European Union (DiDAX, 101115134). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. Giulia Cavicchioni is member of the INdAM Research Group GNSAGA. Violetta Weger is supported by the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement no. 899987.

### v3

@@ Acknowledgments
Jessica Bariffi was partly supported by the Federal Ministry of Education and Research of Germany in the program of "Souverän. Digital. Vernetzt." Joint project 6G-RIC, project identification number: 16KISK022. She is now funded by the European Union (DiDAX, 101115134). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. Giulia Cavicchioni is member of the INdAM Research Group GNSAGA. She conducted the majority of this research at the University of Trento, Trento, Italy, and is currently affiliated with the Institute of Communication and Navigation, German Aerospace Center, Oberpfaffenhofen, Germany. Violetta Weger is supported by the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement no. 899987.